Assuming that men’s weights are normally distributed with a mean of 172 lb. and a population standard deviation of 29 lb. Find the probability that 36 randomly selected men have a mean weight of less than 167 lb.
Answer #1

Solution :

Given that ,

mean = \mu = 172

standard deviation = \sigma = 29

\sigma\bar x = \sigma / \sqrt n = 29 / \sqrt 36 = 4.8333

P(\bar x < 167) = P((\bar x - \mu \bar x ) / \sigma \bar x < (167 - 172) /4.8333 )   

P(z < -1.03)

0.1515

Probability = 0.1515